I need some help.
Find each of the following for the functions defined by f(x)=x^2-4 and g(x_=x+2
A. (fg)(x)
B. (f/g)(x)

Answer by ewatrrr(24785)

A. (fg)(x) = (x+2)^2 - 4 = x^2 + 4x
B. (f/g)(x)= x-2